QUESTION 18 Find the value of k that makes the antidifferentation formula true (4x - 7) * dx = k(4x – 7)12+C
Answer #1

we know that finan= unt+C Here Given that I= f(ux-7) dhe put un-7= 0 l so asahau / I= .1 + 611 de -- Iz 191 197C I = Go C un